a car of mass 1250 kg and a bus of mass 5000 kg are moving with eqaul linear momentum. find the ratio of their kinetic energies


1
Expert's answer
2020-12-23T10:25:42-0500

By definition, the linear momentum of a body is:


"p = mv"

where "m" is the mass of a body, and "v" is its velocity. Since both cars have equal linear momentum, we can write:


"p_1 = p_2\\
m_1v_1 = m_2v_2"

where "m_1 = 1250kg, m_2 = 5000kg" are the masses, and "v_1, v_2" are the velocities of the first and second cars respectively. Thus, obtain the ratio:


"\dfrac{v_1}{v_2} = \dfrac{m_2}{m_1}\\
\space \\
\dfrac{v_1^2}{v_2^2} = \dfrac{m_2^2}{m_1^2}"

On the other hand, their kinetic energies are given as follows:


"K_1 = \dfrac{m_1v_1^2}{2}\\
K_2 = \dfrac{m_2v_2^2}{2}"

Their ratio is:


"\dfrac{K_1}{K_2} = \dfrac{m_1v_1^2}{m_2v_2^2}"

Substituting the expression for "v_1^2\/v_2^2", obtain:


"\dfrac{K_1}{K_2} = \dfrac{m_1\cdot m_2^2}{m_2\cdot m_1^2} = \dfrac{m_2}{m_1}\\
\dfrac{K_1}{K_2} =\dfrac{5000kg}{1250kg} = 4"

Answer. 4.